# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

### v1

@@ 3.2. Analysis for Rule 1
Finally, by taking into consideration the entire conditional, we arrive at
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\mathbb{E}[\|y^{\delta}-F(u^{\delta}_{k})\|^{2}]$
This with (3.7), (3.5) for $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
By incorporating Jensen’s inequality (2.2) with $f(x)=x^{2}$, we conclude that
This means that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. Further, by engaging the triangular inequality (2.1) and the fact that $\|u_{0}-\hat{u}\|\leq\rho$, we conclude that
$\mathbb{E}[\|u^{\delta}_{k+1}-u_{0}\|]$ $\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]+\mathbb{E}[\|\hat{u}-u_{0}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$.
Next, we will prove that $\hat{k}_{\delta}$ is finite. To establish this, we can utilize 3.2 with (3.13) and (3.7) along with the fact that $1-\lambda_{k}\leq 1$ for $\lambda\in[0,\frac{1}{2})$ to get
$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
((3.15)) $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}+2\lambda_{k}\rho c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.\hskip 56.9055pt$
Using the inequality that $\lambda_{k}^{2}\leq\lambda_{k}$ and the definition of the measurability of $u_{k}^{\delta}$ along with the additivity of expected values with respect to $\mathcal{F}_{k}$ in 3.2 to get that
((3.16)) $\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+D\|y^{\delta}-F(u_{k}^{\delta})\|^{2}$ $\leq\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}$
$\hskip 56.9055pt+2\rho c(\rho)\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
Taking the sum of (3.16) from $k=0$ to $k=\hat{k}_{\delta}-1$ to deduce that
((3.17)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\right).$
From (3.5) and inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})\hskip 5.69054pt\forall\hskip 5.69054pta,b\in\mathbb{R}$, we get
((3.18)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\left(\|F(u_{k}^{\delta})-y^{\delta}\|+\frac{M}{k+a}\right)^{2}\leq 2\left(\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+\frac{M^{2}}{(k+a)^{2}}\right).$
By taking the sum of (3.18) from $k=0$ to $\hat{k}_{\delta}-1$, we get
((3.19)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}\leq 2\left(\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right).$
Using (3.17) in (3.19) to get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}$ $\leq 2\left[D^{-1}\left\{\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}\right.\right.$
$\left.\left.\hskip 85.35826pt+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right\}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right].$
By applying (3.8) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$, we further get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}<\infty.$
Hence $\hat{k}_{\delta}$ is finite. Finally, from the definition of $\hat{k}_{\delta}$ in (3.5), we get
$\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as\quad}\delta\to 0.$
Therefore, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. ∎
@@
Let Assumption 3.1 hold and let the sequence $\{\lambda_{k}\}$ be chosen such that (3.7) and (3.8) hold. Additionally, let $\{y^{\delta}\}$ represent a family of noisy data that fulfill Assumption 3.2. Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the integer determined by Rule 1. Then
$\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0\quad\text{as}\quad\delta\to 0.$
Consequently, we have
$\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\to 0\quad\text{and}\quad k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0\quad\text{as}\quad\delta\to 0.$
@@
((3.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}C_{1},$
@@
$\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
@@
((3.21)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\leq D^{-1}C_{1}.$
@@
[…]
@@
$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\Psi(k_{*}(y^{\delta}),y^{\delta})\leq D^{-1}C_{1}.$
Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Further, we note that
$a\|F(u^{\delta}_{k^{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
@@
Hence, we conclude that $\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\to 0$ and $k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0$ as $\delta\to 0,$ since $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Thus, result. ∎
The next result demonstrates that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that (b) of Lemma 3.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
@@
Let all the assumptions of Lemma 3.2 hold. Additionally, assume that
((3.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{k})(1-3\eta)+\omega^{2}L^{2}\geq 0.$
Then, if $u^{\delta}_{k}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k_{*}$, it follows that $u^{\delta}_{k+1}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0})$, provided $\delta$ is sufficiently small.
@@
For $k=0$, result is trivial. By assuming $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k^{*},$ we show that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. From 3.2 and the inequalities $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ and $\|u_{0}-\hat{u}\|\leq\rho$, it follows that
@@
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2})$
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
@@
Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda\in[0,\frac{1}{2})$, we get
@@
Further, the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$ implies that
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}-\|u_{k}^{\delta}\ -\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1-\lambda_{k})(1+\eta)\|y^{\dagger}-y^{\delta}\|^{2}$
$\hskip 28.45274pt+\left(\Omega(1-\lambda_{k})(1-3\eta)-\omega^{2}L^{2}-\kappa^{2})\right)\|F(u_{k}^{\delta})-y^{\delta}\|^{2}.$
By considering the conditional in its whole and utilizing $1-\lambda_{k}\leq 1$, we further reach at
@@
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$

### v3

@@ 2.2. Analysis for Heuristic Rule 1
This with (2.6), $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
[…]
$\|u^{\delta}_{k+1}-u_{0}\|$ $\leq\|u^{\delta}_{k+1}-\hat{u}\|+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$. The desired result is thus established. ∎
@@
We first show that Algorithm 2 is well-defined. Specifically, for
$\|y_{i}^{\delta}-y_{i}^{\dagger}\|_{i\in\{0,1,\ldots,P-1\}}>0,$
the modified discrepancy principle ((2.4)) yields a finite stopping index, ensuring that the algorithm terminates in finitely many steps. Moreover, we prove that $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. Here, the expectation $\mathbb{E}$ and probability $\mathbb{P}$ are taken with respect to the filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$, where $\mathcal{F}_{k}$ is the $\sigma$-algebra generated by the random indices $i_{n}$ for $0\leq n<k$. The formal result is stated below.
Theorem 2.1.
Under the assumptions of Lemma 2.1, the stopping index $\hat{k}_{\delta}$ as defined in (2.4) almost surely remains finite for $\|y_{i}^{\delta}-y_{i}^{\dagger}\|_{i\in\{0,1,\ldots,P-1\}}>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
Proof.
Our objective is to establish that
$\mathbb{P}(\hat{k}_{\delta}=\infty)=0\quad\text{for }\|y_{i}^{\delta}-y_{i}^{\dagger}\|>0,\quad i=0,1,\ldots,P-1.$
By the definition of $\hat{k}_{\delta}$, the event $\Psi:=\{\hat{k}_{\delta}=\infty\}$ can be equivalently expressed as
$\Psi=\left\{\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|+\frac{M}{(k+a)}>\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|;i=0,1,\ldots P-1;\quad\text{for all integer }k\geq 0\right\}.$
We now establish the desired result. Recall from (2.2), ((2.12)), and ((2.6)) that for $\lambda_{k}\in[0,\tfrac{1}{2})$ and $\omega_{k}\in[\omega,\Omega]$, we have
$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}\big{(}1+\kappa^{-2}\omega_{k}^{2}L^{2}\big{)}\rho^{2}$
$\quad+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\,\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
Since $(1-\lambda_{k})^{2}\leq 1$ and $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ by Lemma 2.1, the above yields
$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}\big{(}1+\kappa^{-2}\omega_{k}^{2}L^{2}\big{)}\rho^{2}$
((2.14)) $\quad+2\lambda_{k}\rho\,c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
Next, using $\lambda_{k}^{2}\leq\lambda_{k}$ and $\omega_{k}\leq\Omega$, and noting that $u_{k}^{\delta}$ is $\mathcal{F}_{k}$–measurable while $i_{k}$ is chosen uniformly from $\{0,1,\dots,P-1\}$ independent of $\mathcal{F}_{k}$, we take conditional expectation with respect to $\mathcal{F}_{k}$. By the independence of $i_{k}$, the expected residual term becomes the average over all $P$ components. Thus (2.2) implies
$\mathbb{E}\big{[}\|u_{k+1}^{\delta}-\hat{u}\|^{2}\,\big{|}\,\mathcal{F}_{k}\big{]}+\frac{D}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}$
((2.15)) $\quad\leq\mathbb{E}\big{[}\|u_{k}^{\delta}-\hat{u}\|^{2}\,\big{|}\,\mathcal{F}_{k}\big{]}+\rho^{2}\big{(}1+\kappa^{-2}\Omega^{2}L^{2}\big{)}\lambda_{k}+2\rho\,c(\rho)\,\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
We now take full expectation in (2.2) and sum from $k=0$ to t1t-1. The left–hand side telescopes in the first term, while the residual terms remain additive. Using $\mathbb{E}[\|u_{t}^{\delta}-\hat{u}\|^{2}]\geq 0$, we arrive at
((2.16)) $\begin{split}\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]&\leq\frac{1}{D}\bigg{(}\mathbb{E}[\|u_{0}^{\delta}-\hat{u}\|^{2}]\\ &\quad+\big{(}\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho\,c(\rho)\big{)}\sum_{k=0}^{t-1}\lambda_{k}\\ &\quad+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{t-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\bigg{)}.\end{split}$
Finally, recalling that the initial guess $u_{0}^{\delta}$ is deterministic (or $\mathcal{F}_{0}$–measurable) and satisfies $\|u_{0}^{\delta}-\hat{u}\|\leq\rho$. Hence $\mathbb{E}[\|u_{0}^{\delta}-\hat{u}\|^{2}]=\|u_{0}^{\delta}-\hat{u}\|^{2}\leq\rho^{2}$. So the bound ((2.16)) simplifies to
((2.17)) $\begin{split}\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]&\leq\frac{1}{D}\bigg{(}\rho^{2}+\big{(}\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho\,c(\rho)\big{)}\sum_{k=0}^{t-1}\lambda_{k}\\ &\quad+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{t-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\bigg{)}.\end{split}$
By applying (2.7) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$ in ((2.17)), we further get
((2.18)) $\sum_{k=0}^{t-1}\mathbb{E}\left[\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}\right]=\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]<\infty.$
Now, define $\chi_{\Psi}$ as the characteristic function associated with the set $\Psi$, such that
$\chi_{\Psi}(w)=\begin{cases}1,&\text{if }w\in\Psi,\\ 0,&\text{if }w\notin\Psi.\end{cases}$
By definition of $\Psi$, for every integer $k\geq 0$ we have, for all $i=0,1,\dots,P-1$,
$\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|>\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|-\frac{M}{k+a}.$
Using the elementary inequality $(l-m)^{2}\geq\tfrac{1}{2}l^{2}-m^{2}$ with $l=\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|$ and $m=\dfrac{M}{k+a}$, we obtain for each $i$ on $\Psi$
$\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}\geq\tfrac{1}{2}\tau^{2}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}-\frac{M^{2}}{(k+a)^{2}}.$
Multiplying by the indicator $\chi_{\Psi}$ and taking expectation, then averaging over the uniformly sampled index $i_{k}$, gives
$\begin{split}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}&\geq\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\chi_{\Psi}\big{]}\\ &=\mathbb{E}\Big{[}\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}\chi_{\Psi}\Big{]}\\ &\geq\frac{\tau^{2}}{2P}\,\mathbb{E}\Big{[}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\chi_{\Psi}\Big{]}-\frac{M^{2}}{(k+a)^{2}}\,\mathbb{E}[\chi_{\Psi}]\\ &\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,\mathbb{P}(\Psi)-\frac{M^{2}}{(k+a)^{2}}\,\mathbb{P}(\Psi).\end{split}$
Sum the last inequality over $k=0,\dots,t-1$ and use linearity of expectation, we arrive at
((2.19)) $\sum_{k=0}^{t-1}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}t\,\mathbb{P}(\Psi)-M^{2}\,\mathbb{P}(\Psi)\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}.$
By ((2.18)) there exists a constant $\mathfrak{C}>0$ independent of $t$ such that
$\sum_{k=0}^{t-1}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}\leq\mathfrak{C}.$
Combining this with ((2.19)) gives
$\mathfrak{C}\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t\,\mathbb{P}(\Psi)-M^{2}\,\mathbb{P}(\Psi)\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}.$
Rearrange to obtain
$\mathbb{P}(\Psi)\left(\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t-M^{2}\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}\right)\leq\mathfrak{C}.$
The series $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$, hence
$\mathbb{P}(\Psi)\left(\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t-\frac{M^{2}\pi^{2}}{6}\right)\leq\mathfrak{C}.$
Letting $t\to\infty$, we observe that the term in parentheses on the left–hand side grows without bound whenever $\mathbb{P}(\Psi)>0$. This would contradict the boundedness of the left–hand side unless $\mathbb{P}(\Psi)=0$. Therefore, we conclude that
$\mathbb{P}(\Psi)=0,\quad\text{i.e.,}\quad\mathbb{P}(\hat{k}_{\delta}=\infty)=0.$
This implies that the complement of the event $\Psi$ occurs with probability one, i.e., along almost every sample path, ((2.4)) holds for some finite $\hat{k}_{\delta}$. Finally, from the definition of $\hat{k}_{\delta}$ in (2.4), we get
$\frac{M}{(\hat{k}_{\delta}+a)}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as \quad}\delta\to 0.$
Consequently, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. This completes the proof. ∎
The proof of the following lemma draws inspiration from the methodologies outlined in [36] and [32]. In this result, we examine the asymptotic behavior of the stopping index prescribed by Rule 1 when applied to noisy data satisfying Assumption 2.2. It is noteworthy that the proof of this lemma relies essentially on Theorem 2.1.
@@
Let the sequence $\{\lambda_{k}\}$, $\omega$ and $\Omega$ be chosen such that (2.6) and (2.7) hold and let Assumption 2.1 hold. Additionally, let $\{y^{\delta}\}$ represent a set of noisy data that fulfill Assumption 2.2. Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the stopping index as defined in Rule 1. Then
$\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0\quad\text{as}\quad\delta\to 0.$
Thus, we have
$\mathbb{E}\left[\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\right]\to 0\quad\text{and}\quad\mathbb{E}\left[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\right]\to 0\quad\text{as}\quad\delta\to 0.$
@@
((2.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\mathbb{E}\left[\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]\leq PD^{-1}C_{1},$
@@
$\sum_{i=0}^{P-1}\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
@@
((2.21)) $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\mathbb{E}\left[\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]\leq PD^{-1}C_{1}.$
@@
$\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{k+a}\geq\sum_{k=0}^{\hat{k}_{\delta}-1}\int_{k}^{k+1}\frac{1}{t+a}dt=\int_{0}^{\hat{k}_{\delta}}\frac{1}{t+a}dt=\log\frac{\hat{k}_{\delta}+a}{a}.$
@@
$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\leq PD^{-1}C_{1}.$
Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0$ as $\delta\to 0$. Further, we note that
$a\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k_{*}(y^{\delta})})-y_{i}^{\delta}\|^{2}=a\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
@@
Hence, we conclude that $\mathbb{E}\left[\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\right]\to 0$ and $\mathbb{E}\left[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\right]\to 0$ as $\delta\to 0,$ since $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0$ as $\delta\to 0$. Thus, result. ∎
The next result demonstrates that the iterates $u_{k}^{\delta}$ generated by Algorithm 1 remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that Lemma 2.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
@@
Let all the conditions of Lemma 2.2 be satisfied. Additionally, assume that
((2.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{\text{max}})(1-3\eta)+\omega^{2}L^{2}>0.$
[…]
@@
[…]
@@
$\quad+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
@@
$\quad+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda_{k}\in[0,\frac{1}{2})$, we get
@@
By taking the total conditional and utilizing $1-\lambda_{k}\leq 1$, we further reach at
@@
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)\mathbb{E}[k_{*}\|y^{\delta}-y^{\dagger}\|^{2}].$